OpenStudy (anonymous):

Properties of logarithms help? Simplify 7 log3k + 6 log3m -9 log3n

OpenStudy (anonymous):

choices:

OpenStudy (anonymous):

@lovelyr0se

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

@jdoe0001 @e.mccormick

OpenStudy (mathmale):

I'm going to write out your expression using Equation Editor for greater clarity: \[7\log _{3}k + 6\log _{3}m-9\log _{3}n\]

OpenStudy (mathmale):

Now remember: there are three rules of logs that may apply here. (suggest you write them down for later reference). log a + log b = log (ab) log a - log b = log (a/b) log a^b = b log a Apply the third rule to \[7\log _{3}k:\] Move that " 7 " so as to make it the exponent of k: \[7\log _{3}k=\log _{3}k ^{7}\]

OpenStudy (mathmale):

Do the same thing with the remaining two terms. Is this enuf help to enable you to complete the problem solution? When you're done, you should have only ONE logarithm.